OpenStudy (anonymous):

how would I find x in the equation: 2x(x) = 50 ?

OpenStudy (maheshmeghwal9):

@ipacpc

OpenStudy (maheshmeghwal9):

is ur question this?\[2x \times x=50.\]

OpenStudy (anonymous):

yes

OpenStudy (maheshmeghwal9):

then;\[2x(x)=50.\]\[2x^2=50.\]\[x^2=\frac{50}{2}=25.\]\[x^2=25.\]\[x=\sqrt{25}=\sqrt{5^2.}\]\[x=\pm5.\]

OpenStudy (maheshmeghwal9):

any doubt???????

OpenStudy (maheshmeghwal9):

yes or no?

OpenStudy (maheshmeghwal9):

sorry there should be\[\sqrt{25}=\sqrt{(\pm5)^2}.\]

OpenStudy (anonymous):

Oh, I get it now. Thanks

OpenStudy (maheshmeghwal9):

yw:)